OpenStudy (anonymous):

If cos2(theta) =1, a value of theta is...

OpenStudy (shamim):

\[\cos2\theta=1 \]

OpenStudy (shamim):

\[or, \cos 2\theta=\cos 0\]

OpenStudy (shamim):

\[or, 2\theta=0\]

OpenStudy (shamim):

\[or, \theta=0\]

OpenStudy (anonymous):

sorry i dont understand, 0 isn't one of the choices

OpenStudy (shamim):

ok

OpenStudy (shamim):

trying again

zepdrix (zepdrix):

Well then you should list the choices for us, because 0 is indeed a possibility.

OpenStudy (shamim):

show me choices

OpenStudy (anonymous):

45,90,180,270

OpenStudy (shamim):

\[\cos 2\theta=1\]

OpenStudy (anonymous):

those are all in degrees

OpenStudy (shamim):

\[or, \cos 2\theta=\cos 360\]

OpenStudy (shamim):

\[or, 2\theta=360\]

OpenStudy (shamim):

\[or, \theta=180\]

OpenStudy (anonymous):

so 0 and 180 are the only possibilities?

OpenStudy (shamim):

no

OpenStudy (shamim):

many more possibilities

OpenStudy (shamim):

360 degree is another possib

OpenStudy (shamim):

its infinite number of possibilities

OpenStudy (shamim):

thanks

OpenStudy (anonymous):

sorry i was trying to understand it, had to scroll up a few times, was going to thank you lol, so Thanks :)

OpenStudy (shamim):

i m going to bathroom. coming soon

OpenStudy (shamim):

did u understand my solution

OpenStudy (anonymous):

cos2(theta)=360 cos(theta)=180

OpenStudy (shamim):

no

OpenStudy (shamim):

its not the right way

OpenStudy (anonymous):

please explain

OpenStudy (shamim):

sorry

OpenStudy (shamim):

it was my mistake

OpenStudy (shamim):

i wanna solve it again

OpenStudy (anonymous):

ok, take your time

OpenStudy (shamim):

\[\cos2\theta=\cos360\]

OpenStudy (shamim):

\[2\theta=360\]

OpenStudy (shamim):

\[or, \theta=180\]

OpenStudy (shamim):

do u need any more help

OpenStudy (anonymous):

i understand now, thankyou!